OpenStudy (anonymous):

Regina typically runs at a rate of 4.2 miles per hour. If she ran at this rate for 48 minutes, how far did she run? Use the distance formula, d = rt

OpenStudy (anonymous):

hmm, i think it would be 4.2 times (46/60) = 3.22 miles

OpenStudy (anonymous):

3.36 mi, 5.3 mi. 33.6 mi. or 201.6 mi.

OpenStudy (anonymous):

4.2mi/h * (1h/60min)=0.07 mi/min .07*48=3.36miles

OpenStudy (anonymous):

my bad, 48* haha

OpenStudy (anonymous):

wow